Question 2 The market for barley has the following demand and supply schedule. Price (RM) 100 200 300 400 500 600 Quantity demanded (tonnes) 7000 6000 5000 4000 3000 2000 (tonnes) 1000 2000 3000 4000 5000 6000 (a) State the price, quantity and revenue received at equilibrium. 18 marks] Study shows that barley is good for diabetics. The quantity demanded is increased by 2000 tonnes. Determine the new quantity demanded, price (i) and quantity. 6 marks] (b) Calculate the new revenue. Did the revenue increase or decrease? 6 marks]
Answer #1

Answer:

  1. Equilibrium point: is the point where quantity demanded is equal to the quantity supplied. According the given data equilibrium is at 400 price level where quantity demanded and quantity supplied is equal to 4000 tonnes of barley. We can also identify the equilibrium through the graph as shown in the below graph:

560 400 2o o lo o 4 2 3 4 5 6 7 Quantity (housand Toonas

X- axis shows quantity demand and supplied of barley in thousand tonnes

Y-axis shows the price QD is the quantity demanded of barley

QS is the quantity supplied

The equilibrium price and quantity will be decided at the point where both curve intersect each other

E is the equilibrium point

Equilibrium price is = 400

Equilibrium quantity is = 4000 tonnes

Revenue= quantity * price

Revenue = 4000 * 400 = 16,00,000

So the revenue received at equilibrium is 16,00,000

answer: (i)

If the quantity demanded has increased by 2000 tonnes the demand and supply schedule will be as follows :

Price (RM)

Quantity demanded

Quantity supplied

100

9000

1000

200

8000

2000

300

7000

3000

400

6000

4000

500

5000

5000

600

4000

6000

The quantity demanded has increased so the equilibrium price, quantity and the revenue received at new equilibrium will also change.

The new quantity demanded and quantity supplied are equal at 500 price of barley where the demand and supply both are equal to 5000 tonnes.

The equilibrium quantity is = 5000 tonnes of barley

The equilibrium price is = 500

The revenue received is = 5000 * 500

The revenue received at equilibrium is = 25,00,000

   

Answer: (b)

the new revenue is 25,00,000

The revenue has increased because with the increase in quantity demanded of barley the price of barley also increases and to meet the increased demand of barley the producers also increase the supply of barley